OpenStudy (anonymous):

If the vertex of graph is (5,-4) and two point of the graph are (-7,0) & (-3,0)...what is the discriminant?

OpenStudy (anonymous):

do you know what discriminant is at all?

OpenStudy (anonymous):

what does it show?

OpenStudy (anonymous):

we can recover the whole equation from this information. and you have a choice of methods. since the vertex is \[(5,-4)\] it must be \[y=a(x-5)^2-4\] and now you can replace x by 7, set the result = 0 and solve for a

OpenStudy (anonymous):

it only shows the vertex and two points of the parabola

OpenStudy (anonymous):

or you can say it must look like \[y=a(x-7)(x+3)\] and solve that way

OpenStudy (anonymous):

use foil?

OpenStudy (anonymous):

also you can have \[y=ax^2+bx+c\] and plug in all three points and solve system of equation

OpenStudy (anonymous):

actually you have too much information. i will solve one way, maybe tomas another

OpenStudy (anonymous):

vertex is (5,-4) so it must be \[y=a(x-5)^2-4\] now you know that if x = 7, y = 0 so put \[0=a(7-5)^2-4\] \[a(4)-4=0\] \[4a=4\] \[a=1\] so your equation is \[y=(x-5)^2-4\]

OpenStudy (anonymous):

or since you know \[a=1\] and it must be \[y=a(x-7)(x+3)\] then it is \[y=(x-7)(x-3)\] either way.

OpenStudy (anonymous):

notice in the first example i only used two points. this is all you need and this problem is ill posed. you can recover the equation from the vertex and one zero

OpenStudy (anonymous):

to find the disc, it would be 25+16

OpenStudy (anonymous):

so its 41?

OpenStudy (anonymous):

if you multiply out you get \[y=x^2-4 x-21\]

OpenStudy (anonymous):

then compute \[b^2-4ac=(-4)^2-4\times 1\times (-21)\]

OpenStudy (anonymous):

i get 100

OpenStudy (anonymous):

because the zeros are rational numbers, the discriminant must be a perfect square

OpenStudy (anonymous):

oh thats not one of the choices.. the choices are A. 16 B. 0 C. -13 D. 15

OpenStudy (anonymous):

ah that is because i made a mistake. i thought you said the zeros were (7,0) and (-3,0)

OpenStudy (anonymous):

i missed the minus sign in front of the 7

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

so lets try again. it must be \[y=a(x-5)^2-4\] and you know if x = -3 you get 0

OpenStudy (anonymous):

ok i think it is just \[y=(x+3)(x+7)\] and so it is \[y=x^2+10 x+21\] and therefore \[b^2-4ac=16\]

OpenStudy (anonymous):

ok..thanks for showing the steps..